OpenStudy (lilmike234):

The x-intercept of the graph of f(x)=3log(x+5)-2 is ?

OpenStudy (freckles):

to find x-intercept set y=0 so replace f(x) with 0 and solve for x 0=3log(x+5)-2

OpenStudy (lilmike234):

And then I'll get 10^2/3+5 right ?

OpenStudy (freckles):

so you add 2 on both sides then divided 3 on both sides then do 10^ on both sides then subtract 5 on both sides? \[2=3 \log(x+5) \\ \frac{2}{3}=\log(x+5) \\ 10^\frac{2}{3}=x+5\]

OpenStudy (freckles):

it looks you added 5 on one side and subtracted 5 on the other remember you are suppose to do the same thing to both sides so if you subtract 5 on that one side then you need to subtract 5 on the other